OpenStudy (anonymous):

HELP PLEASE !!!!!!! Solve. -3/4q=12 q =

OpenStudy (anonymous):

-3 --- q=12 4

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@Photon336

OpenStudy (anonymous):

@rishavraj

OpenStudy (photon336):

\[\frac{ -3 }{ -3 }q = \frac{ 48 }{ -3 } = -16\]@kidknowsnothing The first thing we want to do is isolate q we notice we have 3/4 so what we can do is multiply both sides by 4 \[-\frac{ 3 }{ 4 }q = 12\] \[-\frac{ 3 }{ 4 }*\frac{ 4 }{ 1 } = 12*4\] \[-3q = 48\] divide both sides by -3 \[\frac{ -3 }{ -3}q =\frac{ 48 }{ -3 } = -16\] this means that q = -16 we can check this so that we know our answer is correct. \[-\frac{ 3 }{ 4 }*-16 = \frac{ 48 }{ 4 } = 12\]